# Can Charge Transfer Across C–H···O Hydrogen Bonds Stabilize Oil Droplets in Water?

### v1

@@ 2.3 Surface charge and oil-water emulsion stability
Using the non-linear Poisson Boltzmann equation34, 35, 36, 37, we consider how different estimates of surface charge give rise to changes in surface potential which is a measure of emulsion stability. The relationship between the surface potential $\psi_{0}$ and the surface charge density $\sigma_{0}$ can be expressed as:
@@
where $\varepsilon_{r}$ is the relative permittivity, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, $e$ is the elementary charge, and $R$ is the droplet radius. We note that previous experiments have reported average surface droplet radii of 100 nm15, 8 to 230 nm38, 18; here we use a radius of 125 nm as reported most recently18, 8. We also consider the relationship between surface charge and surface potential under different ionic concentrations, where Z is the ion valence and $\rho_{\text{bulk}}$ is the bulk ion density. While a concentration of hydroxide ions of 0.0001 mM/L represents the theoretical ionic strength of pure water, potential measurements are typically conducted in the presence of electrolytes and hence at higher ionic strengths.
@@
Figure 3A shows that the surface charge density of -0.0007 e-/nm2 found from ALMO-EDA, which defines a true quantum mechanical definition of CT, would predict an unstable emulsion at any value of ionic strength, as the surface potential would be below --30 mV. Even under the assumption of a high surface charge density of -0.015 e-/nm2 the emulsions would be unstable at moderately high pH or salt strength, in contradiction to their experiments (i.e. oil droplet in 1mM NaOH or NaCl bulk solution)18. Oil-water emulsions are in fact stable over weeks to months8 at higher pH values, so CT cannot be the origin of oil charging as we have shown it is negligibly small.
[…]
@@
Figure 3: Estimates of surface charge and zeta potentials. (A) Surface potential as a function of surface charge density ($\sigma_{0}$) for a 100 nm-radius spherical interface at various ionic strengths (1 mM, 0.01 mM, 0.0001 mM) using the non-linear Poisson Boltzmann equation34, 35, 36. Dashed vertical lines are surface charge densities of -0.015 17, -0.002 14, and -0.0007 e-/nm2 from ALMO-EDA. See Supplementary Information. (B). An illustration of the potential at the surface as assumed by a CT mechanism, and due to the double layer or beyond under the assumption of ionic species such as hydroxide ions.Adapted with permission from41. Copyright 2025 American Chemical Society
@@
Pullanchery et al.18 also reported a simulation in which aqueous neopentane moved in the presence of an electric field. However, we re-analyzed the uploaded trajectories from their publication, finding that the observed net drift primarily corresponded to the center-of-mass motion of the entire simulated system, but not neopentane itself. This is because the oil carries no significant net charge from CT, and it thus explains why there is a lack of net drift under an applied electric field in the AIMD simulations of neopentane in water (see Supplementary Figure S2 and discussion).18
@@ 3 Discussion and Conclusion
Quantifying CT is rigorously possible using computational quantum chemistry calculations, which provide direct access to the electron density and even via the approximate wavefunction. Our results from ALMO-EDA19, 20, which defines a true quantum mechanical definition of CT as electron relaxation from an optimally polarized state to a delocalized state, allows for both forward and backward CT contributions to lowering the system’s energy. In the water dimer, CT is predominantly unidirectional, with the hydrogen bond acceptor oxygen transferring charge to the donor hydrogen, while backward CT is almost negligible. In contrast, forward CT in the water–hexane complex occurs from the hydrogen bond acceptor oxygen to the donor hydrogen in hexane, but this is largely balanced by backward CT from hexane to the hydrogen of water. This bidirectional nature of CT weakens the directionality of hydrogen bonding, contributing to the more flexible and distorted hydrogen-bonding geometry observed at the water–oil interface, and most relevant here, yields a negligible surface charge value of --0.0007 e-/nm2 that accumulates from oil to water. As we have shown, such insignificant amounts of surface charge would result in an instability of the oil-water emulsion, in contradiction to experimental observations.
@@
Atomic charge partitioning methods assign electron density to individual atoms, which therefore includes a larger effect of partitioning the charge distribution of the optimally polarized state.19 This leads to an overestimation of the CT value and an unphysical estimate of the effective donor-acceptor energy gap42, 43. Hence the very large surface charge of --0.015 e-/nm2 from charge partitioning schemes used by Hassanali and co-workers are inherently flawed.17, 15 This is evident from the slow and smooth 1/r decay of charge accumulation deep into the oil phase in their theoretical analysis (see Figure 217 and Figure 818), behavior that instead signifies electrostatic screening rather than what is expected for true CT43 which is highly localized at the interface as we have shown above.
@@
Since there is no net drift of the neopentane under an applied electric field, the electric-field-induced time-dependent polarization (also referred to as “dynamic polarization”) of neopentane cannot result in a sustained directional force as reported.18 Although dynamic polarization may give rise to transient forces when a molecule is exposed to an electric field, these effects are inherently short-lived, whereas the electrophoretic mobility reflects a steady-state phenomenon governed by the linear response of charges within the slip plane that defines the zeta potential. Regarding the pH dependence, it has been suggested that pH modifies electronic polarization of the bulk solution which induces mobility differences between hydroxide and chloride ions. However, it is known that Br- and I- both have greater polarizability than Cl- but there is almost no change in the mobility and zeta potential among the ions. A second explanation invokes the Grotthuss mechanism, suggesting that the conductivity of OH- is higher than that of Cl-. Yet the molar conductivity of OH- is lower than that of H3O+44; if mobility scales with conductivity, acidic solutions should exhibit higher mobility not hydroxide.
@@
Furthermore, previous and unambiguous EDA studies have explicitly stated that CT would red shift, not blue shift, the C-H vibrational modes of hydrophobic groups interacting with water.25, 26, 22 Given the experimental observation by Roke and co-workers that their C-H vibrations blue shift, then CT as a mechanism must be ruled out, and instead is attributable to Pauli repulsion that arises as a confinement effect at the interface between oil and water. This is supported by the observation that the ”frozen” term in the ALMO-EDA is a majority of the energetic stabilization from C-H···O hydrogen bonds as we have shown in Figures 1 and 2. This is one of the necessary conditions for the VSFS spectroscopic features to be observed to blue-shift, i.e. relatively weak CT that is insufficient to compensate for the blue-shifting effect of the frozen interaction.22
Since the CT mechanism is clearly unviable as an explanation of emulsion stability and electrophoresis, as well as the VSFS measurements themselves, what is the molecular explanation for oil charge? The net oil charge must be large as it has been shown to give rise to large electric fields of $\sim$ --55 to --90 MV/cm at the interface.8 Assertions that impurities are responsible are also unlikely given the orders of magnitude increase in surface area of an emulsion compared to a planar interface. Furthermore their concentration would need to be large enough to exhibit some charging effect and yet at the same time such concentrations should be visible in the SFG and SCRS experiments, which they are not.15, 8
[…]
The implication of these related studies is that the surface potential and zeta potential are in fact not equivalent as assumed by Roke and co-workers, that ionic species are present and must be accounted for in the interpretation of their VSFS experiments, and clearly they should not be ignored in any physical models or simulations. At the same time the surface affinities of hydroxide and hydronium to simplified planar interfaces using simulation models are way too small to explain the needed surface charge densities to induce mobility and maintain emulsion stability. Hence it is still an open question at present as to the molecular origin that enhances the surface affinity of (presumably) hydroxide, but which no one has yet identified.
4 Methods
Charge Transfer Decomposition Analysis. The ALMO-EDA20 scheme separates the overall intermolecular interaction energy, $\Delta E_{\text{INT}}$, into contributions from frozen interaction ($\Delta E_{\text{FRZ}}$), polarization ($\Delta E_{\text{POL}}$), and charge transfer ($\Delta E_{\text{CT}}$) from the corresponding intermediate states
The frozen state is obtained by directly antisymmetrizing the isolated fragment wavefunctions. The polarization state is defined by the relaxation of the molecular orbitals of each fragment in the presence of other fragments and is CT-free.55, 56 In the large basis set limit, orbital overlap between fragments can cause charge transfer contamination in the polarized state. The second generation of ALMO-EDA resolved this by retaining only fragment electric response functions57, 58(FERFs), ensuring accurate polarized states even at the complete basis set limit. Finally, the CT state is achieved from the SCF calculation of the whole complex without any constraint, and it includes the electron flow to and from each molecule in the system due to intermolecular relaxation or mixing of the molecular orbitals.19, 29 The net direction of electron flow is determined by the balance between forward and backward contributions, processes that contribute to stabilizing the system by lowering its energy.
In the ALMO-EDA framework the amount of CT is defined as
$\Delta Q=\text{Tr}\{\hat{P}_{\text{CT}}\}-\text{Tr}\{\hat{P}_{\text{POL}}\hat{P}_{\text{CT}}\},$ (4)
where $\hat{P}_{\text{POL}}$ and $\hat{P}_{\text{CT}}$ are the density operators of the POL and CT states. $\Delta Q$ is the electron count in the final state that lies outside the density operator of the initial state, and is identical59 to the excitation number.60 The CT decomposition analysis of ALMO-EDA breaks this quantity into fragment additive pairs. For instance, the amount of charge transfer from fragment $x$ to fragment $y$ is obtained as
$\Delta Q_{x\rightarrow y}=\sum_{ia}\langle\psi^{ya}|\hat{P}^{\text{eff}}_{vo}|\psi_{xi}\rangle\langle\psi^{xi}|\hat{X}_{ov}|\psi_{ya}\rangle,$ (5)
where $\hat{P}^{\text{eff}}_{vo}$ is an effective density operator, $\hat{X}_{ov}$ is the generator of the unitary transformation connecting $\hat{P}_{\text{POL}}$ and $\hat{P}_{\text{CT}}$, and the $\psi$’s are the covariant and contravariant ALMOs. We can rotate the occupied and virtual ALMOs such that $\hat{X}_{ov}$ is diagonal under the rotated orbitals. These special orbitals form a most compact basis for describing the charge transfer, and they are called the complementary occupied-virtual pairs (COVP). Typically, the CT between two fragments in the CT process can be described by one or two dominant COVPs.
Given the balance between computational cost and accuracy, we selected def2-TZVPD as the optimal basis set for EDA (Details in Supplementary Figure S3). We note on technical grounds the basis set used by Poli and co-workers did not include diffuse functions17, although diffuse functions are necessary for properly describing electron density delocalization and hydrogen bond interactions, and their absence leads to misrepresenting the extent of CT in hydrogen-bonded networks61.
Due to the substantial size of the interfacial hexane–water clusters, which contain on average $\sim$400 atoms per cluster, performing EDA at the $\omega$B97X-V/def2-TZVPD level of theory poses a significant computational challenge, particularly for the evaluation of the polarized state. To alleviate the computational cost associated with polarization, we employed two key strategies to enhance efficiency. First, each hexane molecule was assigned as an independent fragment, thereby avoiding the overhead introduced by large fragments. All water molecules were treated collectively as a single fragment. Second, we utilized the recently developed uncoupled-FERFs (uFERFs)58, which provides an enormous speedup through a diagonal approximation of the orbital hessian matrix. The previous benchmark results 58 demonstrated that the polarization energies of neutral systems are almost unaffected by uFERF approximation.
Molecular dynamics simulations The system, consisting of 1450 water molecules and 200 hexane molecules, was simulated using the AMOEBA polarizable force field within the OpenMM62 platform and all calculations were done in double precision. The dimensions of the simulation box were set to 3.5 × 3.5 × 7.5 nm3. Simulations were kept at a constant temperature of 298.15 K and a constant Pzz of 1 atm, where Pzz is the component of the pressure tensor perpendicular to the interface (only the z-dimension of the periodic simulation box is allowed to change while the x- and y-dimensions are kept fixed). The simulations used the MTSLangevinIntegrator for efficient temperature control and integration. Non-bonded interactions were handled using the Particle Mesh Ewald (PME) method, with a non-bonded cutoff of 1.2 nm and an Ewald error tolerance of 0.0005. The system was equilibrated for 1 ns, with a time step of 0.5 fs. Following equilibration, a 20 ns production run was conducted to investigate the structural properties of the water-hexane interface. Snapshots were saved every 100 ps, yielding a total of 200 frames for EDA analysis.
[…]
DATA AVAILABILITY
@@ 4 Methods
ACKNOWLEDGMENT. We thank the CPIMS program, Office of Basic Energy Sciences, Chemical Sciences Division of the U.S. Department of Energy under Contract DE-AC02-05CH11231. MHG and HS acknowledge support from the U.S. National Science Foundation through Grant CHE-2313791. This work used computational resources provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract DE-AC02-05CH11231, and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory (Supported by the Director, Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231).

### v3

@@ 2.3 Surface charge and oil-water emulsion metastability
Using the non-linear Poisson Boltzmann equationGouy1910, Chapman1913, Debye1923, ohshima1982accurate, we consider how different estimates of surface charge give rise to changes in surface potential which is a measure of emulsion stability. For ionic species it is appropriate to use the non-linear Poisson Boltzmann equationGouy1910, Chapman1913, Debye1923, ohshima1982accurate, which connects the surface charge density $\sigma_{0}$ to the surface potential $\psi_{0}$:
@@
where $\varepsilon_{r}$ is the static dielectric constant of bulk water, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, $e$ is the elementary charge, and $R$ is the droplet radius. We note that previous experiments have reported average surface droplet radii of 100 nmpullanchery2021charge, Shi2025 to 230 nmpullanchery2020stability, Pullanchery2024; here we use a radius of 125 nm as reported most recentlyPullanchery2024, Shi2025. We also consider the relationship between surface charge and surface potential under different ionic concentrations, where Z is the ion valence and $\rho_{\text{bulk}}$ is the bulk ion density. While a concentration of OH- ions of 0.0001 mM/L is the ionic strength of pure water, surface potential and electrokinetic measurements are also conducted in the presence of electrolytes and hence at higher ionic strengths that also must be reconciled for any mechanism.
@@
Figure 3A shows that the surface charge density of -0.0007 e-/nm2 found from ALMO-EDA, which defines a true quantum mechanical definition of CT, would predict an unstable emulsion at any value of ionic strength, as the surface potential would be below -30 mV. Even under the assumption of a high surface charge density of -0.015 e-/nm2 suggested by Poli et al.,poli2020charge the emulsions would be unstable at moderately high pH or salt strength, also in contradiction to experiments (i.e. an oil droplet in 1mM NaOH or NaCl bulk solution as reportedPullanchery2024). Oil-water emulsions are in fact metastable over weeks to monthsShi2025 at at all stated pH values, so surface charge is required to be larger as often advocated by Beattie and co-workers.beattie2009surface
But because CT is short-ranged, the surface charge density is a very thin double layer such that the macroscopic Poisson-Boltzmann model is not warranted. Instead we use the integral form of the one-dimensional Poisson equationbecker2022
$\psi(z)=-\frac{1}{\varepsilon_{0}\varepsilon_{r}}\int_{z_{1}}^{z}\left[\int_{z_{1}}^{z^{\prime}}\rho_{e}(z^{\prime\prime})\,dz^{\prime\prime}\right]dz^{\prime}$ (3)
to integrate over the Gaussian distributed charge over the narrow interface to describe the double layer under the CT assumption.The Gaussian-distributed charge density is generated based on an approximation from the surface charge density, representing the spatial distribution of charges along the surface normal. This is illustrated for the -0.015 $e-/nm{{}^{2}}$ value in Figure 3B (red dashed line) which is very similar to the reported AIMD simulations by Poli et alpoli2020charge. As also seen in Figure 3B the resulting surface potential (solid lines) is very small and unconducive to metastability, and more realistic estimates of surface charge from ALMO-EDA simply reinforces this conclusion. Even changing the bulk dielectric of 80 to the high frequency limit of 1.8 under the assumption that dipolar fluctuations are completely damped out at the interface would still result in surface potentials well below -30 mV.
@@
Figure 3: Estimates of surface charge and zeta potentials. (A) Surface potential as a function of surface charge density ($\sigma_{0}$) for a 100 nm-radius spherical interface at various ionic strengths (1 mM, 0.01 mM, 0.0001 mM) using the non-linear Poisson Boltzmann equationGouy1910, Chapman1913, Debye1923. Dashed vertical lines are surface charge densities of -0.015 poli2020charge, -0.002 vacha2012charge, and -0.0007 e-/nm2 from ALMO-EDA. See Supplementary Information.(B) The surface potential (solid lines) as a function of surface charge density using the Poisson equation.
(C). An illustration of the potential at the surface as assumed by a CT mechanism, and due to the double layer or beyond under the assumption of ionic species such as hydroxide ions. Adapted with permission fromlacour2025role. Copyright 2025 American Chemical Society.
Electrophoresis experiments performed on oil-water emulsions measure the electrokinetic mobility of the oil droplet, $\mu_{e}$ and its connection to the $\zeta$ potential to explain emulsion metastability (Figure 3C). The lateral electric field applied in an electrokinetic experiment must couple to an electric currentBonthuis2010, for which mobile charges are needed in order to distinguish them from bound charges, and to justify their use in the Helmholtz-Smoluchowski equation
$\mu_{e}=\varepsilon_{0}\varepsilon_{r}\zeta/\eta$ (4)
For ionic species, using the bulk water dielectric constant and viscosity in Eq. 4 yields good agreement with electrokinetic experiments for oil-water emulsions. Under the CT assumption, however, it is clear that Eq. 4 is invalid for a number of reasons. First is because it requires the localized CT electron density to be divided into immobile and mobile regions to create a slip plane which undergoes shear to generate mobility.lacour2025role Another concern is that at an interface with immobile charges one would have to use a value closer to the high-frequency water dielectric constant in Eq. (4) since water orientational and vibrational contributions have been damped out, and again reducing mobility by a factor of 20 to 40 that is not in line with experiments.
@@
It is also possible that the CT density on the oil is fully bound and negligibly small, and thus does not produce a finite conductivity, and hence there is no $\zeta$ potential or mobility at all. In fact, Pullanchery et al.Pullanchery2024 also reported an AIMD simulation in which a neopentane molecule in water moved in the presence of an applied electric field. However, we analyzed their uploaded AIMD trajectories from their publication, finding that the observed net drift corresponded to the center-of-mass motion of the entire simulated box, but not neopentane itself. This is because the oil carries no significant net charge from CT, and it thus explains why there is a lack of net drift under an applied electric field in their AIMD simulations (see Supplementary Figure S2 and discussion).Pullanchery2024
Finally, the absence of net drift indicates that the electric-field-induced time-dependent polarization (also referred to as “dynamic polarization”) of neopentane alone cannot produce sustained directional forces as reported.Pullanchery2024 Although dynamic polarization may give rise to transient forces when a molecule is exposed to an electric field, these effects are inherently short-lived, whereas the electrophoretic mobility reflects a steady-state phenomenon governed by the linear response of charges within the slip plane that defines the zeta potential. Since no drift is observed in their DFT simulations, consistent with other AIMD simulations of the air-water interface where CT is operative that also found no field-induced flowbecker2022, the CT mechanism nor dynamic polarization produces a finite conductivity that can give rise to a $\zeta$ potential.
@@ 3 Discussion and Conclusion
Quantifying CT is rigorously possible using computational quantum chemistry calculations, which provide direct access to not only the electron density but the wavefunction and a molecular orbital basis to describe quantum mechanical CT. Our results from ALMO-EDAkhaliullin2008analysis, horn2016probing, defined as electron relaxation from an optimally polarized state to a delocalized state, allows for both forward and backward CT contributions to lowering the system’s energy. In the water dimer, CT is predominantly unidirectional, with the hydrogen bond acceptor oxygen transferring charge to the donor hydrogen, while backward CT is almost negligible. In contrast, forward CT in the water–hexane complex occurs from the hydrogen bond acceptor oxygen to the donor hydrogen in hexane, but this is largely balanced by backward CT from hexane to the hydrogen of water. This bidirectional nature of CT weakens the directionality of hydrogen bonding, contributing to the more flexible and distorted hydrogen-bonding geometry observed at the water–oil interface, and most relevant here, yields a negligible surface charge value of -0.0007 e-/nm2 that accumulates from oil to water. As we have shown, such insignificant amounts of surface charge would result in an instability of the oil-water emulsion, in contradiction to experimental observations.
@@
Atomic charge partitioning methods assign electron density to individual atoms, which therefore includes a larger effect of partitioning the charge distribution of the fully polarized state.khaliullin2008analysis This leads to an overestimation of the CT value and an unphysical estimate of the effective donor-acceptor energy gapkhaliullin2009electron, khaliullin2013microscopic. Hence the very large surface charge of -0.015 e-/nm2 from charge partitioning schemes used by Hassanali and co-workers are inherently flawed.poli2020charge, pullanchery2021charge This is evident from the slow and smooth 1/r decay of charge accumulation deep into the oil phase in their theoretical analysis (see Figure 2poli2020charge and Figure 8Pullanchery2024), behavior that instead signifies electrostatic screening rather than what is expected for true CTkhaliullin2013microscopic which is highly localized at the interface as we have shown above.
@@
Beyond quantifying the magnitude of charge transfer, another question is whether such interfacial electron density can be shear-separated by a lateral electric field to produce steady-state electrokinetic flow. This would require the transferred charge on one molecule to hop onto neighboring molecules, enabling lateral current flow under an applied field.Bonthuis2010 However, the DFT-based AIMD simulations reported by Poli et al. exhibit no net drift for neopentane in waterpoli2020charge, and is consistent with other AIMD simulations which have shown no field-induced flow.becker2022 Regarding the pH dependence, it has been suggested that pH modifies electronic polarization of the bulk solution which induces mobility differences between OH- and Cl- ions. However, it is known that Br- and I- both have a greater polarizability than Cl- but there is almost no change in the mobility and zeta potential among the ions. A second explanation invokes the Grotthuss mechanism, suggesting that the conductivity of OH- is higher than that of Cl-. Yet the molar conductivity of OH- is lower than that of H3O+rodriguez2015exact; if mobility scales with conductivity, acidic solutions should exhibit higher mobility compared to OH-.
@@
Furthermore, previous and unambiguous EDA studies have explicitly stated that CT would red shift, not blue shift, the C-H vibrational modes of hydrophobic groups interacting with water.gu1999fundamental, li2002physical, Mao2019 Given the experimental observation by Roke and co-workers that their C-H vibrations blue shift, then CT as a mechanism must be ruled out, and instead is attributable to Pauli repulsion that arises as a confinement effect at the interface between oil and water. This is supported by the observation that the ”frozen” term in the ALMO-EDA is a majority of the energetic stabilization from C-H···O hydrogen bonds as we have shown in Figures 1 and 2. This is one of the necessary conditions for the VSFS spectroscopic features to be observed to blue-shift, i.e. relatively weak CT that is insufficient to compensate for the blue-shifting effect of the frozen interaction.Mao2019
Since the CT mechanism is clearly unviable as an explanation of emulsion stability and electrophoresis, as well as the VSFS measurements themselves, what is the molecular explanation for oil charge? The net oil charge must be large as it has been shown to give rise to large electric fields of $\sim$ -55 to -90 MV/cm at the interface.Shi2025 In our view it is still an open question at present as to the molecular origin that enhances the surface affinity of negatively charged species to the oil-water and related hydrophobic interfaces, and we must consider all available evidence regarding the remaining hypotheses of ionic charge from water dissociationBeattie2004, mccarty2008electrostatic, gray2009explanation, beattie2012, beattie2014a, Gan2017, pullanchery2021charge, Djerdjev2025 or arising from surface impuritiesRoger2012, uematsu2019impurity, uematsu2020nanomolar. Both explanations face serious challenges and yet neither should be dismissed as the explanation at this stage. We hope to provide a clarifying review on this topic in the near future.
DATA AVAILABILITY. All data is available upon request from the authors.
@@
ACKNOWLEDGMENT. We thank the U.S. National Science Foundation through Grant CHE-2313791 for support of the EDA work and analysis. We thank the CPIMS program, Office of Basic Energy Sciences, Chemical Sciences Division of the U.S. Department of Energy under Contract DE-AC02-05CH11231 supporting research on the interfacial oil-water emulsions. This work used computational resources provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract DE-AC02-05CH11231, and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory (Supported by the Director, Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231).